A balloon of gross weight W newton is descending down with an acceleration of f unit. The weight that must be removed or dropped down so as to give the balloon an equal upward acceleration.
  1. Wf/g
  2. 2Wf/g
  3. 2Wf/g+f
  4. W(g+f)/f
The problem I face is formation of the equations.

Grade:10

1 Answers

Eshan
5 years ago
Dear student,

The mass of the balloon, if its weight is W is [m=\dfrac{W}{g}]

The bouyant force [B] acting the body is same independent of the direction of movement of balloon. Hence while moving down,

W-B=mf

While moving up,

B-W'=mf

Hence the weight that must be removed=W-W'=2mf=2\dfrac{W}{g}f